OpenStudy (anonymous):

8 times the square of a nonzero number is equal to 96 times the number. What is the number?

OpenStudy (anonymous):

ok first write out the euation: 8*x^2=96*x

OpenStudy (anonymous):

Alrighty.....

OpenStudy (anonymous):

then try to solve it as a normal equation... x^2=96-8*x

OpenStudy (anonymous):

x^2=82*x

OpenStudy (anonymous):

88?

OpenStudy (anonymous):

sorry i did the math wrong its x^2=88*x

OpenStudy (anonymous):

then you move over the x to the x^2 side by dividing it, whats x^2 divided by x?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

It's 12. I got it

OpenStudy (anonymous):

yep,actualy the answer is 88

OpenStudy (anonymous):

8x^2=96x 8x^2-96x=0 8x(x-12)=0 8x=0 -------- x-12=0 x=12

OpenStudy (anonymous):

12 and 88 are both answers :)